A school anniversary committee is to be selected from all its employee. There are 21 male where 16 are faculty members and 5 are non-teaching personnel, and 39 are female where 24 are teachers and 15 are non teaching personnel. What is the probability that the first person selected is female?
1
Expert's answer
2020-04-30T19:07:50-0400

Total employees = 21 +39 =60

female = 39


"P(first person =female) =\frac{39}{60}=0.65"